I need your help how do you solve x^2-2x-9=0

Answer by mathgeek102(3)

x^2-2x-9=0
1.) You would subtract 2x from it's side to the other side of the equation (where zero is)
2.) Do the same with the -9
3.) You should end up with x^2=-2x-9
2.) Next you divide x from both sides
4.) This should end up as x=-2-9 which equals -11
Hope I helped